An article in the ASCE Journal of Energy Engineering (Overview of Reservoir Release Improvements at 20 TVA Dams, Vol. 125,
Answer #1

a)

sample mean 'x̄= 3.231
sample size    n= 20
std deviation s= 2.1290
std error ='sx=s/√n=2.129/√20= 0.4761
for 95% CI; and 19 df, value of t= 2.093 from excel: t.inv(0.975,19)
margin of error E=t*std error    = 0.996
lower bound=sample mean-E = 2.235
Upper bound=sample mean+E = 4.227
from above 95% confidence interval for population mean =(2.23<μ<4.23)

b)

std errror of mean ='sx=s*√(1+1/n)= 2.182
for 95% CI; and 19 df, value of t= 2.093 from excel: t.inv(0.975,19)
margin of error E=t*std error    = 4.57
lower bound=sample mean-E = -1.335
Upper bound=sample mean+E= 7.797
from above 95% prediction interval for population mean =(-1.34 ,7.80)

c)

for sample size n=20& 95% tolerance level with 99% CI value of k= 3.1680
sample mean                  x̅= 3.231
standard deviation       σ= 2.129
margin of error              E =    k*std deviation 6.745
lower tolerance level     = sample mean -E                                                = -3.514
Higher tolerance level   = sample mean +E                                                = 9.976

95% tolerance interval =(-3.51 , 9.98)